ethan needs to determine the number and type of roots for the equation using the discriminant. his solution is shown.explain what ethan did correctly and what he did incorrectly and then correct mistakes.
1. b^2-4ac=-5^2-4(1)(10)=-25-40=-65
2. since -65<0, there are 2 non real solutions.

Answer by stanbon(75887)

ethan needs to determine the number and type of roots for the equation using the discriminant. his solution is shown.explain what ethan did correctly and what he did incorrectly and then correct mistakes.
1. b^2-4ac=-5^2-4(1)(10)=-25-40=-65
2. since -65<0, there are 2 non real solutions.
-------
Mistake: b^2 is NOT -5^2 = -25
--------------------
Correct: b^2 = (-5)^2 = 25
===============
Cheers,
Stan H.
===============